OpenStudy (gabylovesyou):

@ganeshie8 (5 - 4i)(3 + 3i)

OpenStudy (trojanpoem):

Normal multiplication

OpenStudy (trojanpoem):

i * i = i^2 which is rootof(-1) ^2 = -1

OpenStudy (gabylovesyou):

i get 15 + 15i - 12i - 12i ?

OpenStudy (gabylovesyou):

whoops 12i^2 at the end

OpenStudy (trojanpoem):

+12

OpenStudy (gabylovesyou):

but -4 * 3 is -12

OpenStudy (trojanpoem):

It is.

OpenStudy (trojanpoem):

I was talking about 12 i ^2 = 12

Nnesha (nnesha):

 when you multiply 3i times -4i you should get -12i^2 right ?